### v2

# SCREP: Scene Coordinate Regression and Evidential Learning-based Perception-Aware Trajectory Generation

Abstract
Autonomous flight in GPS‑denied indoor spaces requires trajectories that keep visual‑localization error tightly bounded across varied missions. Map‑based visual localization methods such as feature matching require computationally intensive map reconstruction and have feature‑storage scalability issues, especially for large environments. Scene coordinate regression (SCR) provides an efficient learning-based alternative that directly predicts 3D coordinates for every pixel, enabling absolute pose estimation with significant potential for onboard robotics applications. We present a perception‑aware trajectory planner that couples an evidential learning‑based SCR pose estimator with a receding-horizon trajectory optimizer. The optimizer steers the onboard camera toward reliable scene coordinates with low uncertainty, while a fixed-lag smoother fuses the low-rate SCR pose estimates with high-rate IMU data to provide a high-quality, high-rate pose estimate. In simulation, our planner reduces translation and rotation RMSE by at least 4.9% and 30.8% relative to baselines, respectively. Hardware‑in‑the‑loop experiments validate the feasibility of our proposed trajectory planner under close-to-real deployment conditions.
I Introduction
Vision-based localization has made rapid progress in recent years, contributing to advances in unmanned aerial vehicle (UAV) navigation systems for GPS-denied indoor environments. Map-based visual localization approaches leverage prior environmental knowledge to provide drift-free pose estimates, making trajectory planning that considers localization performance crucial for reliable navigation in such environments. Most perception-aware trajectory planning approaches [13, 39, 30, 4, 34, 40, 24, 12, 11, 38, 23] steer the camera toward feature-rich regions to improve the pose estimate obtained by structure from motion (SfM) or visual-inertial odometry (VIO). However, SfM-based feature maps require computationally intensive reconstruction and have feature‑storage scalability issues, especially for large environments, making them challenging for onboard robotics applications. Moreover, even with perception-aware guidance, VIO-based methods are still limited by cumulative error.
(a) stata1
(b) stata2
Figure 1: Top‑down 3‑D reconstructions of the two sites produced by SCR. (left) RGB-colored points (right) Normalized entropy map. The accumulated scene coordinates and their uncertainties from multiple viewpoints reveal spatially consistent uncertainty distributions across the environment.
Figure 2: Illustration of the proposed receding-horizon perception‑aware trajectory generation. (blue) Camera orientations are optimized to steer the camera toward low-entropy coordinates. (red) Position trajectory (black) Current pose.
Scene coordinate regression (SCR) offers a learning-based alternative that directly predicts a 3D coordinate for every pixel from a single image [33, 6, 7, 27, 37, 5, 36, 17]. A Perspective‑n-Point (PnP) solver then produces a drift‑free absolute camera pose, achieving high metric accuracy even over long trajectories while requiring only a compact neural network. However, not all pixels produce reliable coordinates. Moreover, SCR often produces outliers and runs at low frequency with high latency, which complicates its use in real-time planning and control.
This paper proposes a perception-aware trajectory planning algorithm that explicitly utilizes the pixel-wise uncertainty of an evidential learning-based SCR network (E-SCRNet). To the best of our knowledge, this is the first active-perception work using SCR. We train a neural network for SCR with evidential learning [1], so each pixel outputs scene coordinates and both aleatoric and epistemic uncertainties. We use the entropy as the uncertainty metric since it jointly captures aleatoric and epistemic uncertainties. The entropy map drives a receding-horizon trajectory optimizer that (i) weighs low-entropy pixels higher and (ii) removes high-entropy pixels from the cost, thereby orienting the camera toward informative viewpoints while following the given reference position trajectory. A fixed-lag smoother then fuses the low-rate SCR pose with high-rate IMU measurements, reducing latency, mitigating outlier SCR pose estimates, and eliminating drift caused by IMU bias.
The main contributions of this work are as follows:
  • We present a real‑time, receding-horizon trajectory planner that injects E-SCRNet uncertainty into perception-aware trajectory optimization, facilitating high-accuracy localization.
  • We empirically show that entropy is a more informative uncertainty metric than separate aleatoric and epistemic uncertainty measures.
  • We validate our approach through simulation and hardware‑in‑the‑loop experiments, where the proposed planner outperforms baselines in localization.
II Related Work
II-A Perception-Aware Trajectory and Path Planning
PAMPC [13] augments nonlinear model predictive control (MPC) with a visibility cost and Zhang et al. [39] frame perception–aware planning in a receding-horizon fashion. Yaw optimization was later coupled with differential flatness and a differentiable camera cost [30], extended to incorporate semantic priors [4], and parameterized time‑optimally [34]. Other directions incorporate information-theoretic [40] or topological metrics [24].
More recent work combines position–yaw optimization with a novel feature‑visibility model [11], considers feature‑limited or unknown environments [38], introduces control barrier functions for safety guarantees [23], or is more data-driven [12]. However, all of these approaches rely on feature matching (FM) or visual–inertial odometry (VIO). Unlike other methods, our work addresses perception-aware planning under scene coordinate regression (SCR).
II-B Learning-based Visual Localization
Beyond FM and VIO, which utilize information from visual features, numerous learning-based visual localization methods have emerged. Absolute Pose Regressors (APRs) employ neural networks to directly regress camera pose [22, 20, 35]. Uncertainty‑aware variants [19, 29] also estimate uncertainty, yet APRs remain markedly less accurate than both model‑based pipelines and other learning‑based alternatives [32].
HLoc [31] retrieves candidate keyframes with global descriptors and refines the pose via local feature matching. While accurate, its memory and computation grow almost linearly with the number of keyframes and can exceed the resources available on onboard hardware as the map grows.
SCR regresses a dense 3‑D scene‑coordinate map and then estimates the pose with PnP-RANSAC. The evolution of SCR began with random forests [33], which DSAC later replaced with differentiable RANSAC and CNNs [6, 7] for end-to-end training. Without depth supervision, ACE [5] enables SCR to be trained in a few minutes. Recent advances include hierarchical SCR [27, 37] and a pretrained image encoder for co-visibility among images [36, 17] to handle large scenes, with [17] additionally improving robustness. Although SCR attains high absolute accuracy with a single forward pass, it can produce outliers. Also, its application in robotics remains underexplored.
II-C Uncertainty Quantification
Bayesian neural networks (BNNs) estimate aleatoric and epistemic uncertainty via Monte‑Carlo dropout [14], deep ensembles [26], or more general Bayesian learning frameworks [21]. Uncertainty‑aware APR [19] and VIO using BNNs and neural radiance fields [16] show the benefit of incorporating BNN uncertainty estimates for localization. However, BNNs require multiple forward passes for the uncertainty quantification.
On the other hand, evidential learning offers closed‑form aleatoric and epistemic estimates without sampling [1, 9] and has been adopted for NeRF‑aided SCR [10], semantic segmentation [2], semantic mapping [25], and traversability learning [8]. By applying evidential learning to our approach, the uncertainty of each predicted scene coordinate is quantified with only a single forward pass.
Figure 3: System architecture. (1) A pose estimation block consisting of E‑SCRNet, PnP‑RANSAC, and a fixed‑lag smoother; (2) A two‑stage trajectory optimization block for position and yaw.
III System Overview
Given generic navigation scenarios, this study aims to improve an initial positional path by simultaneously optimizing the vehicle’s position and yaw trajectories.
Problem. At each step, we aim to compute a path of length $T_{\mathrm{plan}}$ in a receding-horizon fashion based on the following inputs: (i) reference position waypoints or trajectory, (ii) predicted scene coordinates and their associated uncertainties from E-SCRNet, and (iii) the current UAV pose estimate. The objective is to produce a dynamically feasible trajectory that (a) follows the reference, (b) respects velocity/acceleration limits, and (c) orients viewpoints with low uncertainty to maximize localization accuracy. Thanks to the differential flatness of multirotor dynamics [28], the position and yaw trajectories $\{\mathbf{p}(t),\psi(t)\}$, and their derivatives fully determine the required state and control.
Pipeline. Fig. 3 depicts the pipeline for our receding-horizon trajectory generation framework. Images from a camera pass through E‑SCRNet [10] to produce scene coordinates and uncertainties, which a PnP‑RANSAC module converts into camera poses. To compensate for the latency and outliers of this stream, we fuse it with high‑rate IMU measurements using a fixed‑lag smoother, yielding a clean, high‑frequency state estimate.
This pose estimate, the most recent scene coordinates and uncertainties from the previous execution period $T_{\text{exec}}$, and reference waypoints feed a perception‑aware trajectory optimization. This optimization is split into position and yaw subproblems in order to achieve the runtime efficiency required for real‑time execution. The resulting reference trajectory $\{\mathbf{p}^{\star}(t),\psi^{\star}(t)\}$ is continuously generated until the UAV reaches its goal pose.
IV Methodology
In this section, we first describe the training of E‑SCRNet, an SCR network trained via deep evidential regression [1], and discuss its representation of uncertainty in the context of perception-aware trajectory planning. We then present our B‑spline–based receding-horizon perception-aware trajectory planner that utilizes these scene coordinates and their associated uncertainties.
IV-A Scene Coordinate Regression with Deep Evidential Regression
We train a neural network for SCR and assess the uncertainty of each predicted scene coordinate so that the perception‑aware trajectory optimization can reason about perception quality. We adopt deep evidential regression [1], following the approach in [10]. In contrast to BNNs, deep evidential regression can jointly quantify both aleatoric and epistemic uncertainty without requiring multiple forward passes. Unlike previous works [1, 10], our approach represents the total uncertainty as entropy to identify reliable scene coordinates.
Let $\mathbf{v}\in\mathbb{R}^{3}$ denote the scene coordinate of each pixel. For brevity, we write $v$ when referring to one Cartesian component. We model $v\sim\mathcal{N}(\mu,\sigma^{2})$, where the unknown mean and variance $\mathbf{t}=(\mu,\sigma^{2})$ are themselves random variables drawn from a Gaussian and inverse gamma prior $\mu\sim\mathcal{N}(\gamma,\sigma^{2}/\lambda\bigr)$, $\sigma^{2}\sim\Gamma^{-1}(\alpha,\beta)$. Assuming that the priors on $\mu$ and $\sigma^{2}$ are factorized, $p(\mathbf{t}|\mathbf{m})$ is the Normal Inverse-Gamma (NIG) distribution, which is a Gaussian conjugate prior parameterized by $\mathbf{m}=(\gamma,\lambda,\alpha,\beta)$ with $\lambda>0$, $\alpha>1$, and $\beta>0$.
The NIG distribution serves as a higher‑order evidential distribution from which the lower‑order Gaussian likelihood that generates scene coordinates is sampled, allowing the network to express the prediction $\mathbb{E}[\mu]=\gamma$}, the aleatoric uncertainty $\mathbb{E}[\sigma^{2}]=\beta/(\alpha-1)$, and epistemic uncertainty $\text{Var}[\mu]=\beta/\big(\lambda(\alpha-1)\big)$.
Deep evidential regression aims to learn the hyperparameters $\mathbf{m}_{\theta}=(\gamma_{\theta},\lambda_{\theta},\alpha_{\theta},\beta_{\theta})$. We can express the likelihood of the scene coordinate given the hyperparameters $\mathbf{m}_{\theta}$ in closed-form $p(v|\mathbf{m}_{\theta})=\text{St}\big(v;\gamma_{\theta},\beta_{\theta}(1+\lambda_{\theta})/(\lambda_{\theta}\alpha_{\theta}),2\alpha_{\theta}\big)$ where $\text{St}\big(v;\eta,\tau^{2},\chi)$ is a Student-t distribution with location $\eta$, scale $\tau^{2}$, and degree of freedom $\chi$.
The total loss $\mathcal{L}_{\theta}$ minimized during training is the mean of the per‑sample loss $\mathcal{L}_{\theta}^{i}=\mathcal{L}_{\theta,\mathrm{nll}}^{i}+\rho\,\mathcal{L}_{\theta,\mathrm{reg}}^{i}$}. The loss $\mathcal{L}_{\theta,\mathrm{nll}}^{i}$ is chosen as the negative log-likelihood (NLL) of $p(v|\mathbf{m}_{\theta})$ for maximum likelihood estimation (MLE):
$\begin{split}&\mathcal{L}_{\theta,\mathrm{nll}}^{i}=\log\!\left(\Gamma(\alpha_{\theta})/\Gamma\!\left(\alpha_{\theta}+\tfrac{1}{2}\right)\right)\\ &+\tfrac{1}{2}\log\!\left(\pi/\lambda_{\theta}\right)-\alpha_{\theta}\log\!\left(2\beta_{\theta}(1+\lambda_{\theta})\right)\\ &+\bigl(\alpha_{\theta}+\tfrac{1}{2}\bigr)\log\!\left(\lambda_{\theta}(v_{i}-\gamma_{\theta})^{2}+2\beta_{\theta}(1+\lambda_{\theta})\right).\end{split}$ (1)
The loss $\mathcal{L}_{\theta,\mathrm{reg}}$ is designed for regularization:
$\mathcal{L}_{\theta,\mathrm{reg}}^{i}=(2\lambda_{\theta}+\alpha_{\theta})|v_{i}-\gamma_{\theta}|.$ (2)
After training, we can express the predictive entropy $H_{\theta}=\mathbb{E}_{v\sim p(v|\mathbf{m}_{\theta})}[-\log p(v|\mathbf{m}_{\theta})]$ in its closed form
$\begin{split}H_{\theta}&=\!(\alpha_{\theta}+\tfrac{1}{2})(\psi(\alpha_{\theta}+\tfrac{1}{2})-\psi(\alpha_{\theta}))\\ &+\tfrac{1}{2}\log\!\left(2\pi\beta_{\theta}(1+\lambda_{\theta})/\lambda_{\theta}\right)\\ &+\log\!\left(\Gamma(\alpha_{\theta})/\Gamma\!\left(\alpha_{\theta}+\tfrac{1}{2}\right)\right),\end{split}$ (3)
where $\psi(\cdot)$ is the digamma function. We use entropy as the uncertainty representation because it captures both aleatoric and epistemic uncertainty. The detailed justification is given in Section V-B. We refer the reader to [1] for more information on deep evidential regression.
IV-B Trajectory Optimization
We receive the current UAV pose as estimated by the fixed-lag smoother alongside reference waypoints and scene coordinates with their corresponding uncertainties. Given these inputs, position and yaw trajectories are optimized to track waypoints and orient reliable scene coordinates. In this section, we detail the perception‑aware trajectory optimization.
Uniform B-Spline. Both position and yaw are parameterized as clamped, uniform B‑splines of degree $k=3$:
$\mathbf{q}(t)=\sum_{i=0}^{n}\mathbf{q}_{i}N_{i,k}(t)$ (4)
where the knot vector is $\{t_{0},t_{1},\dots,t_{n+k+1}\}$, the control points are $\mathbf{q}_{i}=\{q_{x,i},q_{y,i},q_{z,i},q_{\psi,i}\}\in\mathbb{R}^{4}$, and $N_{i,k}(t)$ denote the B‑spline basis functions. Each knot span has equal duration $\Delta t=t_{i+1}-t_{i}$. For clamping, the first and last $k$ knots are repeated: $t_{0}=\dots=t_{k-1}$ and $t_{n+1}=\dots=t_{n+k+1}$. We fix the first control point to the current UAV pose and the last control point to the final waypoint within the planning horizon $T_{\mathrm{plan}}$. The other control points become decision variables during the optimization.
Optimization Cost. Similar to Bartolomei et al. [4], the total cost function, $C_{\mathrm{total}}$, is represented as the weighted sum of the cost functions for different purposes. To reduce computational load, we optimize the position and yaw separately by minimizing two disjoint sub-objectives rather than the full $C_{total}$:
$\begin{gathered}C_{\mathrm{total}}^{\mathbf{p}}=\lambda_{\mathrm{wp}}C_{wp}^{\mathbf{p}}+\lambda_{\mathrm{eq}}C_{\mathrm{eq}}^{\mathbf{p}}+\lambda_{\mathrm{ie}}C_{\mathrm{ie}}^{\mathbf{p}}+\lambda_{s}C_{s}^{\mathbf{p}}\\ C_{\mathrm{total}}^{\psi}=\lambda_{\mathrm{fov}}C_{\mathrm{fov}}^{\psi}+\lambda_{\mathrm{eq}}C_{\mathrm{eq}}^{\psi}+\lambda_{\mathrm{ie}}C_{\mathrm{ie}}^{\psi}+\lambda_{s}C_{s}^{\psi}.\end{gathered}$ (5)
where the superscripts, $\mathbf{p}$ and $\psi$, indicate that the terms are evaluated on the position and yaw trajectory, respectively. For brevity, we will drop the superscripts in the following. The coefficients $\lambda_{(\cdot)}$ balance five components: $C_{\mathrm{wp}}$ penalizes deviation from the given waypoints, $C_{\mathrm{fov}}$ rewards keeping reliable scene coordinates within the camera’s field-of-view (FOV), $C_{\mathrm{eq}}$ enforces the trajectory starts from the given initial velocity, acceleration, and jerk, $C_{\mathrm{ie}}$ enforces the specified velocity and acceleration bounds, and $C_{s}$ promotes smoothness of the trajectory.
Let $\{\mathbf{w}_{0},\dots,\mathbf{w}_{r-1}\}$ denote all waypoints and let $\{s_{1},\dots,s_{r-2}\}$ be the prescribed arrival times for all but the first and last. The waypoint cost $C_{\mathrm{wp}}$ penalizes the squared Euclidean error between the position trajectory and each desired waypoint:
$C_{\mathrm{wp}}=\sum_{i=1}^{r-2}\,\sum_{o\in\{x,y,z\}}\bigl(q_{o}(s_{i})-w_{o,i}\bigr)^{2}.$ (6)
Our formulation of the FOV cost $C_{\mathrm{fov}}$ is inspired by Murali et al. [30]. Whereas the original cost encourages the camera to track as many covisible features as possible within its frustum, our version instead rewards keeping low‑uncertainty scene coordinates in view regardless of covisibility.
We assume the camera is mounted on the UAV with a known relative pose. Given the trajectories, we can compute the camera position $\mathbf{p}_{C,i}\in\mathbb{R}^{3}$ and orientation $R_{i}^{C}\in\mathrm{SO}(3)$ with respect to the world frame at each time $t=s_{i}$. Let $\mathbf{v}_{j}\in\mathbb{R}^{3}$ be the $j$-th scene‑coordinate point in the world frame; the same point expressed in the camera frame is then $\mathbf{v}^{C}_{i,j}=R_{i}^{C}\bigl(\mathbf{v}_{j}-\mathbf{p}_{C,i}\bigr)$.
Next, we need to build a differentiable indicator function describing whether scene coordinates $\mathbf{v}^{C}_{i,j}$ are within the camera frustum. For this purpose, we consider the displacement vectors from the optical center to each of the four image‑plane corners:
$\begin{split}\mathbf{z}_{TR}&=\big[(w-c_{x})l_{x}\quad(-c_{y})l_{y}\quad f\big]^{\top}\\ \mathbf{z}_{LR}&=\big[(w-c_{x})l_{x}\quad(h-c_{y})l_{y}\quad f\big]^{\top}\\ \mathbf{z}_{TL}&=\big[(-c_{x})l_{x}\quad(-c_{y})l_{y}\quad f\big]^{\top}\\ \mathbf{z}_{LL}&=\big[(-c_{x})l_{x}\quad(h-c_{y})l_{y}\quad f\big]^{\top}.\end{split}$ (7)
Here, $w$ and $h$ are the image width and height in pixels, $(c_{x},c_{y})$ is the principal point in pixels, $l_{x}$ and $l_{y}$ are the physical pixel sizes $(m/\text{pix})$, and $f$ is the focal length in meters. The normal vectors of the four side half planes of the camera frustum are $\mathbf{z}_{\mathrm{TR}}\!\times\!\mathbf{z}_{\mathrm{LR}}$, $\mathbf{z}_{\mathrm{TL}}\!\times\!\mathbf{z}_{\mathrm{TR}}$, $\mathbf{z}_{\mathrm{LL}}\!\times\!\mathbf{z}_{\mathrm{TL}}$, and $\mathbf{z}_{\mathrm{LR}}\!\times\!\mathbf{z}_{\mathrm{LL}}$. Using these normals, we construct a soft indicator vector
$\mathbf{o}_{i,j}=\frac{1}{2}\begin{bmatrix}1+\tanh\!\left((\mathbf{z}_{\mathrm{TR}}\!\times\!\mathbf{z}_{\mathrm{LR}})\!\cdot\!\mathbf{v}^{C}_{i,j}/s\right)\\[2.0pt] 1+\tanh\!\left((\mathbf{z}_{\mathrm{TL}}\!\times\!\mathbf{z}_{\mathrm{TR}})\!\cdot\!\mathbf{v}^{C}_{i,j}/s\right)\\[2.0pt] 1+\tanh\!\left((\mathbf{z}_{\mathrm{LL}}\!\times\!\mathbf{z}_{\mathrm{TL}})\!\cdot\!\mathbf{v}^{C}_{i,j}/s\right)\\[2.0pt] 1+\tanh\!\left((\mathbf{z}_{\mathrm{LR}}\!\times\!\mathbf{z}_{\mathrm{LL}})\!\cdot\!\mathbf{v}^{C}_{i,j}/s\right)\\[2.0pt] 1+\tanh\!\left((\mathbf{v}^{C}_{i,j}-f\mathbf{e}_{3})\!\cdot\!\mathbf{e}_{3}/s\right)\end{bmatrix}$, (8)
where $\mathbf{e}_{3}=[0,0,1]^{\top}$ and $s>0$ are the camera’s forward axis and a smoothing constant, respectively. For the $j$-th scene coordinate at time $t=s_{i}$, the desired scalar indicator is defined as the product of the elements of $\mathbf{o}_{i,j}$:
$F(\mathbf{c}_{i},\mathbf{v}_{j})=\prod_{k=1}^{5}o_{i,j,k}.$ (9)
The scalar indicator approaches 1 when the point lies inside the frustum and 0 otherwise. A larger $s$ yields a smoother transition but a less selective indicator.
The cost is the weighted sum of the indicators over all sampling times and scene coordinates:
$C_{fov}=-\sum_{i=1}^{r-1}\sum_{j=1}^{n_{f}}\exp(-a_{\mathrm{fov}}H_{\theta,j})F(\mathbf{c}_{i},\mathbf{v}_{j}).$ (10)
Here, $n_{f}$ denotes the number of selected scene coordinates. These coordinates are randomly chosen among those with entropy values below a predefined threshold. The weight $\exp(-a_{\mathrm{fov}}H_{\theta,j})$ depends on the entropy $H_{\theta,j}$ from Eq. (3); $a_{\mathrm{fov}}>0$ is constant. Lower entropy, i.e., higher certainty, yields a larger weight, encouraging the yaw trajectory to keep such scene coordinates in view.
Because we solve the optimization repeatedly over a moving horizon $T_{\text{exec}}$, each new trajectory must splice smoothly onto the previous one. To enforce continuity, the initial‑equality cost $C_{\text{eq}}$ penalizes any mismatch between the prescribed initial derivatives and those of the candidate trajectory at $t=t_{0}$:
$C_{\text{eq}}=\|\mathbf{v}(t_{0})-\mathbf{v}_{0}\|_{2}^{2}+\|\mathbf{a}(t_{0})-\mathbf{a}_{0}\|_{2}^{2}+\|\mathbf{j}(t_{0})-\mathbf{j}_{0}\|_{2}^{2}$ (11)
where $\mathbf{v}_{0},\mathbf{a}_{0},\mathbf{j}_{0}\in\mathbb{R}^{4}$ denote the desired initial velocity, acceleration, and jerk for the translation coordinates and yaw. Note that the initial position and yaw automatically match their desired values because a clamped B‑spline is anchored to the first control point.
The inequality cost $C_{\text{ie}}$ enforces that every component of the planned trajectory respects user‑specified velocity and acceleration limits:
$C_{\text{ie}}=\!\!\sum_{i=1}^{r-1}\!\sum_{o\in\{x,y,z,\psi\}}\!\bigl(M_{v}(v_{o}(s_{i}))+M_{a}(a_{o}(s_{i}))\bigr),$ (12)
where $v_{o}(s_{i})$ and $a_{o}(s_{i})$ denote, respectively, the velocity and acceleration of coordinate $o$ at sample time $s_{i}$. The function $M_{u}(u_{o})=\max\big((u_{o}^{2}-u_{\max}^{2})^{2},0\big)$ is applied to both velocity $M_{v}(v_{o})$ and acceleration $M_{a}(a_{o})$.
The smoothness cost $C_{\text{s}}$ promotes a smooth trajectory by enforcing consistency between consecutive elastic bands, $\mathbf{S}_{i+1,i}$ and $\mathbf{S}_{i-1,i}$ with $\mathbf{S}_{i,j}=\mathbf{q}_{i}-\mathbf{q}_{j}$:
$C_{\text{s}}=\sum_{i=1}^{n-1}\bigl\lVert\mathbf{S}_{i+1,i}+\mathbf{S}_{i-1,i}\bigr\rVert_{2}^{2}.$ (13)
TABLE I: Localization accuracy for each translation in cm / rotation in . Columns show results for the four yaw strategies, Forward, Vanilla, FIM, and our proposed method (Ours). Accuracy is reported for raw SCR output (“SCR”) and after fusion with IMU data through the fixed-lag smoother (“IMU+SCR”). Best results are highlighted in bold, and second-best results are highlighted in underline.
[…]
V Experiments
We begin by evaluating our approach in simulation. Our evaluation is based on two key assumptions: (A1) lower uncertainty implies higher scene‑coordinate accuracy, which we validate empirically in Section V-B, and (A2) uncertainty varies smoothly over the trajectory since it is spatially correlated, allowing its use within short planning horizons, as supported by Fig. 1 and the supplementary video. Using only current scene coordinates for short horizons rather than predicting future features is justified by the spatial correlation of the uncertainty.
Our two goals are: (a) to verify that the entropy defined in Eq. (3) provides an appropriate measure of uncertainty, and (b) to demonstrate the effectiveness of our perception‑aware trajectory planner. First, we empirically examine how the proposed entropy compares with conventional aleatoric and epistemic uncertainty measures. We then benchmark our planner against baselines. We also conduct hardware-in-the-loop experiments to validate system-level integration with onboard IMU data and simulated visual input.
Figure 4: Plots showing normalized uncertainty and scene coordinate error of E-SCRNet predictions. (left) Mean and standard deviation of L2 error (right) Density histogram.
Figure 5: Qualitative results of E-SCRNet for a test image. (upper left) Aleatoric Uncertainty (upper middle) Epistemic Uncertainty (upper right) Entropy (lower left) L2 Error of predicted scene coordinates (lower middle) Predicted scene coordinates (lower right) Input image.
(a) winter
(b) ampersand
Figure 6: Comparison of baseline planners (Forward, Vanilla, FIM) with the proposed method on “winter” and “ampersand” trajectories. (left) Time histories of translation and rotation errors of the fixed-lag smoother for Forward (blue), Vanilla (orange), FIM (green), and Ours (red) over the full duration of each trajectory. (right) Top‑down view showing reference trajectory (black dotted), the predicted trajectories (colored), camera orientations (green arrows) for one repetition of each trajectory.
V-A Setup
We run photorealistic simulations using the FlightGoggles simulator on an Intel Core i9-7920X CPU [15]. The simulated camera resolution is fixed at $480\times 480$ pixels, with an FoV of $90^{\circ}$. We train the E‑SCRNet following the architecture in [7] for each of the two environments stata1 and stata2 as shown in Fig. 1.
To highlight the advantages brought by our perception‑aware planner, we deliberately use a small training set captured along a single trajectory (“picasso”) from the Blackbirds dataset [3], trained for only 100k iterations. We set the regularization coefficient $\rho=10^{-2}$. Following [7], we train the network using Adam with a learning rate of $10^{-4}$ and a batch size of 1, completing in 14 minutes on one NVIDIA RTX 6000 Ada GPU.
[…]
For the hardware-in-the-loop experiments, the quadrotor flies within an arena while onboard IMU data is logged and the camera stream is simulated with images rendered from the virtual stata1 scene, as shown in Fig. 7(a). The quadrotor’s pose is tracked using a motion capture system. An external machine equipped with an NVIDIA RTX A5000 GPU processes simulator images and executes SCR inference, smoothing, and optimization, transmitting control commands wirelessly to the UAV.
V-B Uncertainty Representation
We collect scene coordinate predictions and their uncertainty maps from E‑SCRNet while the UAV follows a circular trajectory. Fig. 4 analyzes the relationship between the L2 prediction error and three uncertainty estimates: aleatoric, epistemic, and entropy. We clip the largest 3% of uncertainty values and normalize the remainder to $[0,1]$ for visual clarity.
As shown in Fig. 4, the mean error increases with increasing entropy, whereas its correlation with the others is weaker. Moreover, the error’s standard deviation declines as entropy decreases, while it remains high across the full range of the other measures. This empirical evidence directly validates assumption A1 stated at the beginning of this section. Fig. 4 further reveals that entropy is less skewed toward zero than the other uncertainties, indicating that it provides a more informative and well‑spread metric. Fig. 5 shows qualitative uncertainty and error maps for a representative test image.
Based on these observations, we adopt entropy as our uncertainty metric for the scene coordinate predictions since entropy simultaneously captures both aleatoric and epistemic effects and closely tracks the error.
V-C Simulation Results.
Comparison to baselines. Previous VIO-based perception-aware planners [30, 4] use an FOV cost which directs the camera toward visual features tracked over continuous frames in order to improve VIO accuracy. While this is similar to our proposed method, these approaches crucially do not make use of any uncertainty metric. In contrast, the proposed method leverages the uncertainties from our E-SCRNet to inform selection of reliable scene coordinates. To demonstrate the effectiveness of this additional uncertainty information, we compare our planner with three yaw strategies: First, Forward, which keeps the camera aligned with the velocity vector. Second, Vanilla, which employs the same perception-aware trajectory optimization as our proposed method but without incorporating uncertainty from E-SCRNet. Specifically, Vanilla uses a fixed weight in Eq. (10) of 1.0 and does not filter high-entropy pixels during yaw optimization. This baseline provides behavior equivalent to that from previous strategies [30, 4], while also making use of SCR for fairness in the comparison. Finally, Fisher information matrix (FIM) employs the same perception-aware trajectory optimization but replaces the entropy weight in Eq. (10) with the trace of the FIM introduced in Zhang et al. [40]. We evaluate four trajectories from the Blackbird dataset per site, “clover”, “patrick”, “thrice”, and “winter” in stata1, and “ampersand”, “dice”, “halfmoon”, and “sid” in stata2. Each trajectory is flown multiple times.
Table I reports localization accuracy for both the SCR output through PnP-RANSAC (“SCR”) and the fixed‑lag smoother that fuses IMU and SCR (“IMU+SCR”). Our method consistently outperforms the baselines on “IMU+SCR” localization, reducing translation RMSE by 15.0%, 4.9%, and 23.5% and rotation RMSE by 30.8%, 38.6%, and 41.3% compared to Forward, Vanilla, and FIM, respectively. Mean translation errors are reduced by 9.4%, 5.4%, and 19.7%, while mean rotation errors decrease by 27.3%, 38.5%, and 35.1% relative to the baselines. For “SCR” output, our method achieves the lowest mean errors compared to all baselines. Fig. 6 visualizes the estimated positions and camera orientations for selected trajectories, along with the histogram of translation and rotation, providing a visual representation of the quantitative results presented in Table I.
Consequently, the simulation results validate that our perception-aware trajectory planner improves localization accuracy compared to baseline methods.
Computation time. We break down the runtime of the main modules in our pipeline. The average runtimes for (1) SCR & PnP-RANSAC, (2) the fixed-lag smoothing, and (3) the trajectory optimizations are 62.2ms, 0.8ms, and 28.3ms, respectively. Given the execution periods of the optimization and SCR modules, the overall latency is sufficiently low for real-time implementation.
(a) Physical environment
(b) Translation and rotation error
Figure 7: Hardware-in-the-loop experiment. (7(a)) Drone and spaces for the hardware-in-the-loop experiment. (7(b)) Time histories of translation error (up) and rotation error (down) for Forward and our method.
V-D Hardware-in-the-loop Experiment
The quadrotor follows a rose-petal trajectory, $r=3\cos(2\theta)$ for three laps during the hardware-in-the-loop experiment. Throughout the experiment, we compare the localization performance of the Forward baseline with our method.
Translation RMSE and mean error of the fused pose dropped from 29.7cm and 24.6cm to 22.7cm and 20.0cm, respectively, while rotational RMSE and mean error fell from 6.0 and 5.6 to 2.3 and 2.0. Fig. 7(b) shows the errors over time, demonstrating that the error of our approach remains consistently below that of the Forward throughout a representative lap of the rose petal trajectory. The results confirm the effectiveness and feasibility of the proposed method under close-to-real-world conditions.
VI Conclusion
This paper introduces a receding-horizon, perception‑aware trajectory generation framework using evidential learning-based scene coordinate regression (SCR). Our algorithm consistently orients the camera toward scene coordinates with low uncertainty and produces drift‑free, high‑accuracy poses in real-time. Experiments demonstrate that our framework outperforms baseline methods in localization accuracy and is suitable for hardware-in-the-loop flight tests. Future work will integrate obstacle avoidance and test robustness to appearance changes such as lighting variations and object rearrangements.
ACKNOWLEDGMENTS
We used OpenAI ChatGPT for code design and debugging related to the experiments section, and for polishing and improving the clarity of writing across the entire paper. The authors reviewed, verified, and take full responsibility for all content.